Hello tutor, i really need help with this math problem solving question. Help would be greatly appreciated, thank you so much! Question : A car leaves Edmonton travelling at 40 km/h. A second car leaves Edmonton one hour later and travels at 65 km/h. How far from Edmonton does the second car overtake the first? ( note: this is part algebra, so i have to provide and equation, work and a set up ( table ). I tried three times too do this question and i had no luck, all i know is that the table has car 1 & car 2 & you have too list there speeds, times and distance. I really need to get into math pure next year, thank you so much for the help! > God bless.
Answer by sudhanshu_kmr(1152)


Difference between two cars when second car leaves = 40 km

relative velocity of two cars = 65 - 40 = 25 km/h
required time to overtake = 40/25 hr.

distance traveled by second car in this duration = 65 * 40/25 = 104

so, 104 km from the Edmonton.